# Counterfactual Reasoning for Steerable Pluralistic Value Alignment of Large Language Models

### v1

@@ A.2 Scoring Criteria Calibration
To improve the accuracy of the value attributor and clarify the criteria for multi-level assessment, we introduce a criteria calibration procedure [52]. This process relies on a small human-annotated dataset $\mathcal{D}=(q_{i},r_{i},s_{v_{i}}^{human})_{i=1}^{N}$, where $s_{v_{i}}^{human}$ is the human-assigned importance level of value $v_{i}$ based on the context $(q_{i},r_{i})$. Given the original criteria $T_{ori}$, we first prompt the LLM to infer the importance score $s_{v_{i}}^{infer}$ across all labeled data. Next, we calculate the disagreement between the inferred scores and the human annotations, defined as $\delta_{i}=|s_{v_{i}}^{human}-s_{v_{i}}^{infer}|$. Recognizing the disagreements, we instruct the LLM to refine the original criteria through targeted edits, paraphrasing and augmentation, so that a new version that is most likely to mitigate these disagreements is obtained, denoted as $T_{new}$. With the updated criteria $T_{new}$ replacing $T_{ori}$, the process is repeated iteratively, refining the criteria until the deviation from human annotations is minimized.
@@ A.3 Formal Definition and Implementation of SCM
To implement this reasoning process, we prompt an LLM to aggregate the concepts to generate the final response. Since traditional causal reasoning mainly focuses on tabular data and struggles with natural language, we follow recent studies [45, 46, 48] on causal inference with LLMs. Specifically, the counterfactual reasoning process follows three steps, summarized in Tab. 7.
@@ B.1.1 Value Systems
We leverage the widely used Schwartz Theory of Basic Values [14] to define value profiles for the Touché23-ValueEval dataset. This theory proposes ten universal values that reflect broad human motivations and are organized in a circular continuum to indicate their compatibilities and conflicts. These ten values are grouped into four higher-order meta-types: Openness to Change, Conservation, Self-Enhancement, and Self-Transcendence. Tab. 8 summarizes these ten value dimensions and their motivational definitions.
@@
In the DailyDilemma dataset [33], 301 distinct value dimensions are annotated across scenarios. We select the 50 most frequently occurring value dimensions to define our multi-dimensional value profiles. These values cover diverse ethical, social, and personal motivations, and align well with the broader theoretical space defined by Schwartz. We list them in Tab. 9.
@@ B.1.2 Target Value Profiles
Country-level value profiles. We identify the distinctive value orientations of various countries on Schwartz’s basic human values from the European Social Survey (ESS)  [^1] https://ess.sikt.no/en/datafile/450fa78e-68ab-493f-b169-dbc7ab8ffec2 and other real-world surveys in social science [57]. Taking into account geographic diversity, value variation, and the availability of value-related survey results, we select the value profiles of ten countries as the alignment target for Touché23-ValueEval, including the United States, China, India, South Korea, Singapore, Czech Republic, Russia, Netherlands, the United Kingdom and Germany. In the original value survey data, the importance assigned to each Schwartz’s basic value dimension is a real number $1<d<6$, and we map these importance scores to our scoring scale $[1,5]$ by rounding down the values to the nearest integer.
@@ B.3 Licenses for Existing Assets
  • Touché23-ValueEval  [58]: The dataset is released under the CC-BY-4.0 License and is publicly available at https://huggingface.co/datasets/webis/Touche23-ValueEval.
  • DailyDilemma [33]: The dataset is distributed under the CC-BY-4.0 license License and can be accessed at https://huggingface.co/datasets/kellycyy/daily_dilemmas.
  • PVQ (Portrait Values Questionnaire) [59]: The PVQ is a publicly available academic instrument, widely used for value assessment in cross-cultural research. Researchers are encouraged to cite the original work when using the questionnaire items.
@@ B.4 Implement Details
  • Tree of Thought (ToT) [54]: This method structures the reasoning process as a tree, enabling the model to explore multiple intermediate thoughts before converging on a value-aligned response. Code is available at https://github.com/princeton-nlp/tree-of-thought-llm. In our implementation, we set the tree depth to 2. The first level (breadth = 3) involves generating three candidate responses. The most promising candidate is then selected and further refined in the second level (breadth = 2). Finally, the best-optimized response is chosen as the model’s output.
  • Plan and Solve [55]: This method prompts the model to first generate a structured plan aligned with the target value, and then produce the final response by executing that plan step by step. Specifically, the model is first asked to outline a reasoning trajectory or action path toward the response aligned with the desired value profile. It then sequentially follows this path, performing intermediate reasoning steps before synthesizing the final answer. Code is available at https://github.com/AGI-Edgerunners/Plan-and-Solve-Prompting.
@@
  • CultureLLM [19]: This approach fine-tunes LLMs on culturally grounded datasets derived from the World Value Survey (WVS). We finetune a cultural model separately for each country on both LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct, using 500 augmented samples per country. We follow their code for reproduction, available at https://github.com/Scarelette/CultureLLM.
  • CultureBank [20]: This is a benchmark of culture-aware narratives curated from social media Tiktok and Reddit, categorized by cultural themes like traveling and immigration. We fine-tune cultural LLMs with the Tiktok split on both LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct, following the original method open-sourced at https://github.com/SALT-NLP/CultureBank.
  • CulturePark [18]: This approach uses synthetic data generated via role-based discussions prompted in LLMs. We trained this model separately for each country on both LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct using 500 synthetic samples per country, following their open-sourced code at https://github.com/Scarelette/CulturePark.
  • CultureSPA [56]: This approach generates synthetic pairs contrasting cultural-aware and culturally neutral responses. Then, pairs with obvious differences are used to fine-tune the culturally aligned model. Code is available at https://github.com/shaoyangxu/CultureSPA.
  • VIM [26]: This approach constructs question-answering pairs and argument generation samples based on the training set of Touché23-ValueEval dataset, which reflect value orientations closed to the target value profile. Then, we fine-tuned one model for each country-level value profile on both LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct using the constructed data subset. We follow the open-source code at https://github.com/dongjunKANG/VIM.

### v2

@@ A.2 Scoring Criteria Calibration
To improve the accuracy of the value attributor and clarify the criteria for multi-level assessment, we introduce a criteria calibration procedure [liu2023calibrating]. This process relies on a small human-annotated dataset $\mathcal{D}=(q_{i},r_{i},s_{v_{i}}^{human})_{i=1}^{N}$, where $s_{v_{i}}^{human}$ is the human-assigned importance level of value $v_{i}$ based on the context $(q_{i},r_{i})$. Given the original criteria $T_{ori}$, we first prompt the LLM to infer the importance score $s_{v_{i}}^{infer}$ across all labeled data. Next, we calculate the disagreement between the inferred scores and the human annotations, defined as $\delta_{i}=|s_{v_{i}}^{human}-s_{v_{i}}^{infer}|$. Recognizing the disagreements, we instruct the LLM to refine the original criteria through targeted edits, paraphrasing and augmentation, so that a new version that is most likely to mitigate these disagreements is obtained, denoted as $T_{new}$. With the updated criteria $T_{new}$ replacing $T_{ori}$, the process is repeated iteratively, refining the criteria until the deviation from human annotations is minimized.
@@ A.3 Formal Definition and Implementation of SCM
To implement this reasoning process, we prompt an LLM to aggregate the concepts to generate the final response. Since traditional causal reasoning mainly focuses on tabular data and struggles with natural language, we follow recent studies [bhattacharjee2023towards, bhattacharjee2024zero, feder2023data] on causal inference with LLMs. Specifically, the counterfactual reasoning process follows three steps, summarized in Tab. 7.
@@ B.1.1 Value Systems
We leverage the widely used Schwartz Theory of Basic Values [schwartz2012overview] to define value profiles for the Touché23-ValueEval dataset. This theory proposes ten universal values that reflect broad human motivations and are organized in a circular continuum to indicate their compatibilities and conflicts. These ten values are grouped into four higher-order meta-types: Openness to Change, Conservation, Self-Enhancement, and Self-Transcendence. Tab. 8 summarizes these ten value dimensions and their motivational definitions.
@@
In the DailyDilemma dataset [chiu2024dailydilemmas], 301 distinct value dimensions are annotated across scenarios. We select the 50 most frequently occurring value dimensions to define our multi-dimensional value profiles. These values cover diverse ethical, social, and personal motivations, and align well with the broader theoretical space defined by Schwartz. We list them in Tab. 9.
@@ B.1.2 Target Value Profiles
Country-level value profiles. We identify the distinctive value orientations of various countries on Schwartz’s basic human values from the European Social Survey (ESS)  [^1] https://ess.sikt.no/en/datafile/450fa78e-68ab-493f-b169-dbc7ab8ffec2 and other real-world surveys in social science [ralston2011twenty]. Taking into account geographic diversity, value variation, and the availability of value-related survey results, we select the value profiles of ten countries as the alignment target for Touché23-ValueEval, including the United States, China, India, South Korea, Singapore, Czech Republic, Russia, Netherlands, the United Kingdom and Germany. In the original value survey data, the importance assigned to each Schwartz’s basic value dimension is a real number $1<d<6$, and we map these importance scores to our scoring scale $[1,5]$ by rounding down the values to the nearest integer.
@@ B.3 Licenses for Existing Assets
  • Touché23-ValueEval  [mirzakhmedova2023touch]: The dataset is released under the CC-BY-4.0 License and is publicly available at https://huggingface.co/datasets/webis/Touche23-ValueEval.
  • DailyDilemma [chiu2024dailydilemmas]: The dataset is distributed under the CC-BY-4.0 license License and can be accessed at https://huggingface.co/datasets/kellycyy/daily_dilemmas.
  • PVQ (Portrait Values Questionnaire) [schwartz2001extending]: The PVQ is a publicly available academic instrument, widely used for value assessment in cross-cultural research. Researchers are encouraged to cite the original work when using the questionnaire items.
@@ B.4 Implement Details
  • Tree of Thought (ToT) [yao2023tree]: This method structures the reasoning process as a tree, enabling the model to explore multiple intermediate thoughts before converging on a value-aligned response. Code is available at https://github.com/princeton-nlp/tree-of-thought-llm. In our implementation, we set the tree depth to 2. The first level (breadth = 3) involves generating three candidate responses. The most promising candidate is then selected and further refined in the second level (breadth = 2). Finally, the best-optimized response is chosen as the model’s output.
  • Plan and Solve [wang2023plan]: This method prompts the model to first generate a structured plan aligned with the target value, and then produce the final response by executing that plan step by step. Specifically, the model is first asked to outline a reasoning trajectory or action path toward the response aligned with the desired value profile. It then sequentially follows this path, performing intermediate reasoning steps before synthesizing the final answer. Code is available at https://github.com/AGI-Edgerunners/Plan-and-Solve-Prompting.
@@
  • CultureLLM [li2024culturellm]: This approach fine-tunes LLMs on culturally grounded datasets derived from the World Value Survey (WVS). We finetune a cultural model separately for each country on both LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct, using 500 augmented samples per country. We follow their code for reproduction, available at https://github.com/Scarelette/CultureLLM.
  • CultureBank [shi2024culturebank]: This is a benchmark of culture-aware narratives curated from social media Tiktok and Reddit, categorized by cultural themes like traveling and immigration. We fine-tune cultural LLMs with the Tiktok split on both LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct, following the original method open-sourced at https://github.com/SALT-NLP/CultureBank.
  • CulturePark [li2024culturepark]: This approach uses synthetic data generated via role-based discussions prompted in LLMs. We trained this model separately for each country on both LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct using 500 synthetic samples per country, following their open-sourced code at https://github.com/Scarelette/CulturePark.
  • CultureSPA [xu2024self]: This approach generates synthetic pairs contrasting cultural-aware and culturally neutral responses. Then, pairs with obvious differences are used to fine-tune the culturally aligned model. Code is available at https://github.com/shaoyangxu/CultureSPA.
  • VIM [kang2023values]: This approach constructs question-answering pairs and argument generation samples based on the training set of Touché23-ValueEval dataset, which reflect value orientations closed to the target value profile. Then, we fine-tuned one model for each country-level value profile on both LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct using the constructed data subset. We follow the open-source code at https://github.com/dongjunKANG/VIM.